OpenStudy (anonymous):

if the density of methanol is 0.793kg/L,what is its volume needed for making 2.5L of its 0.25M(molar)solution?please explain with steps.

OpenStudy (vincent-lyon.fr):

2.5 L x 0.25 M = 0.625 mole needed Methanol is CH3OH, its molar mass is 12+3+16+1 = 32 g/mol so mass needed is : \(0.625 \;mol\times32\; g/mol\) = 20 g so volume needed is \(\Large \frac{20 g}{793\: g/L}\) = 25.22 mL